# Adapting Large Language Models for Education: Foundational Capabilities, Potentials, and Challenges

### v2

@@ Abstract
Online education platforms, leveraging the internet to distribute education resources, seek to provide convenient education but often fall short in real-time communication with students. They often struggle to offer personalized education resources due to the challenge of addressing the diverse obstacles students encounter throughout their learning journey. Recently, the emergence of large language models (LLMs), such as ChatGPT, offers the possibility for resolving this issue by comprehending individual requests. Although LLMs have been successful in various fields, creating an LLM-based education system is still challenging for the wide range of educational skills required. This paper reviews the recently emerged LLM researches related to educational capabilities, including mathematics, writing, programming, reasoning, and knowledge-based question answering, with the aim to explore their potential in constructing the next-generation intelligent education system. Based on the current development status, we further outline two approaches for an LLM-based education system: a unified approach and a mixture-of-expert (MoE) approach. Finally, we explore the challenges and future directions, providing new research opportunities and perspectives on adapting LLMs for education.
1 Introduction
@@
Education plays a vital role in shaping individuals’ futures as it forms the foundation for providing people with knowledge, skills, and critical thinking abilities. Conventional education systems heavily rely on teachers for imparting knowledge to students, which place a significant demand on educational resources. However, the advent of online education has substantially lowered the cost of accessing these educational materials. Many people are conveniently acquiring knowledge through online courses and exercises. However, online education faces the issue of not being able to interact with students in real-time like a teacher. Each student inevitably encounters unique challenges during their learning journey. Presently, online education platforms typically offer static course videos and exercises, leaving students unable to ask questions or seek dynamic solutions to their specific issues. Therefore, developing a teaching assistant model capable of addressing students’ individual concerns represents a crucial step forward in advancing online education.
Figure 1: An example of LLM-based education systems integrating multiple abilities to solve student problems.
@@
The emergence of large language models (LLMs) instills optimism in creating an intelligent education system. Since the launch of ChatGPT, LLMs have shown exceptional ability in understanding human knowledge and have been widely applied in various professional fields, including recommendation systems (Lin et al., 2023), healthcare (Liu et al., 2023d), economics (Li et al., 2023), and others. However, the challenge intensifies when constructing an LLM-based education system due to the broad spectrum of knowledge encompassed by education, including mathematics, reasoning, writing, and beyond.
@@
The act of LLMs answering students’ questions can be considered as a process that involves using multiple educational capabilities. For instance, as depicted in Figure 1, the question posed by the student requires the education assistant to possess both mathematical and programming skills simultaneously. Therefore, summarizing the development status of LLMs’ education-related capabilities is meaningful for building the next-generation intelligent education system. In this paper, we study LLMs from the perspective of these education-related capabilities and explore the potential of an LLM-based education system, aiming to provide valuable insights into the areas that need further improvement. We present our primary investigation in Figure 2. Our specific focus centers on five distinct capabilities of LLMs:
@@
  • Mathematics: We primarily summarize the current development of LLMs in solving mathematical problems of varying difficulty and types, encompassing basic arithmetic operations, challenging problems, multi-modal problems, and math proof.
@@
  • Programming: In accordance with human programming conventions, we divide LLMs’ programming process into two stages: code writing and code refinement. We review the researches in this area and summarize the remaining problems for code education.
@@
Besides summarizing the current development status of various foundational abilities related to education, we propose two potential approaches for forming an LLM-based education system. One approach involves training a comprehensive language model with multiple capabilities, while the other is based on an LLM controller in a mixture-of-experts framework. We analyze the implementation difficulties and pros and cons of these two approaches.
LLMs hold vast potential in the field of education, and our work aims to offer insights for building an LLM-based intelligent education system. While Gan et al. (2023) anticipated the prospects of LLMs for education by exploring the various roles they can play in the education process, our work differs. We review the evolution of LLMs from the perspective of education-related capabilities. Additionally, we provide foresight into constructing feasible frameworks for LLM-based education systems. Our work emphasizes a comprehensive understanding of LLMs’ educational competencies and explores frameworks that can effectively integrate these abilities into the educational landscape.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, align=left, minimum width=2.5em, s sep=6pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.6em,font=, where level=2text width=5.3em,font=, where level=3text width=9.7em,font=, [Foundational
@@
Capabilities,text width=6.3em [Mathematics, text width=4em [Basic Arithmetic Problems, text width=8em [Goat (Liu and Low, 2023), GENBERT (Geva et al., 2020),leaf, text width=17.5em] ] [Challenging Mathematics Problems,text width=10.3em [WizardMath (Luo et al., 2023a), MeatMath (Yu et al., 2023),leaf, text width=18.5em] ] [Multimodal Problems,text width=6.3em [Mathvista (Lu et al., 2023), Unigeo (Chen et al., 2022b),leaf, text width=17.3em] ] [Mathematical Proof,text width=6em [LeanDojo (Yang et al., 2023b), LEGO-Prover (Xin et al., 2023),leaf, text width=19.5em] ] ] [Writing,text width=2.5em [Text Summarization,text width=6.3em [BRIO (Liu et al., 2022), PROM (Ma et al., 2023b),leaf, text width=15.4em] ] [Grammatical Error Correction,text width=9.5em [CoEdit (Liu et al., 2022), GrammarGPT (Fan et al., 2023),leaf, text width=18em] ] ] [Programming,text width=4.1em [Code Writing, text width=4.8em [CodeLlama (Rozière et al., 2023), WizardCoder (Luo et al., 2023b),leaf, text width=20em] ] [Code Refinement, text width=6em [Codet (Chen et al., 2022a), SEIDR (Liventsev et al., 2023),leaf, text width=18.5em] ] ] [Reasoning, text width= 3em [Supervised Finetuning,text width=6.3em [CAGE (Rajani et al., 2019), Scratchpad (Nye et al., 2021),leaf, text width=18em] ] [Prompt Engineering,text width=6.3em [CoT (Wei et al., 2022), SelfConsistency (Wang et al., 2022),leaf, text width=18.5em] ] [Hybrid Methodologies,text width=6.3em [FineTuneCoT (Ho et al., 2022), STaR (Zelikman et al., 2022),leaf] ] ] [Knowledge-based QA , text width= 6.3em [Open-domain QA [Retro (Borgeaud et al., 2022), FRESHPROMPT (Vu et al., 2023),leaf, text width=19.8em] ] [Domain-specific QA,text width=6.5em [Reta-llm (Liu et al., 2023c), LeanContext (Arefeen et al., 2023),leaf]] ] ] ] ]
Figure 2: Summary of LLM’s education-related foundational capabilities.
2 Foundational Capabilities
2.1 Mathematics
@@
Mathematics demands reasoning and a sophisticated understanding of complex information, making it one of the disciplines that place the highest premium on these cognitive abilities. There is significant academic interest in enhancing the mathematical capabilities of LLMs (Lu et al., 2022b). In the pursuit of creating an education system based on LLMs, our goal is for it to possess the ability to address various aspects of mathematical problems, whether they involve basic numerical computations, intricate reasoning, or problems integrating multi-modal information. In this section, we summarize the current developments in the mathematical capabilities of LLMs, focusing on four aspects: fundamental numerical computations, complex reasoning, the handling of multi-modal problem-solving, and mathematical proof.
2.1.1 Basic Arithmetic Problems
@@
In human learning mathematics, foundational mathematical operations serve as the basis for addressing more advanced problems. Recently, significant scholarly attention has focused on augmenting LLMs’ proficiency in this domain (Patel et al., 2021; Wang et al., 2017; Zhang et al., 2023a; Geva et al., 2020). Most models try to fine-tune a pre-trained LLM on high-quality mathematics datasets. Given the robust comprehension of human language and notable textual reasoning abilities exhibited by LLMs, it is natural for many to assume that LLMs should effortlessly handle fundamental mathematical problems. However, the reality proves otherwise. Yuan et al. (2023) pointed out that ChatGPT and GPT-4 perform well in addition and subtraction operations, but their accuracy decreases when dealing with multiplication involving larger numerical values. This limitation arises because the language models do not access a calculator during the computations. Yang et al. (2023c) and Liu and Low (2023) proposed to fine-tune LLMs on high-quality datasets and found that even small models could avoid this multi-digit problem. However, it is not feasible to fine-tune each LLM individually. How to prevent such issues during the pre-training phase of language models remains an open question.
2.1.2 Challenging Mathematics Problems

### v3

@@ Abstract.
Online education platforms, leveraging the internet to distribute education resources, seek to provide convenient education but often fall short in real-time communication with students. They often struggle to address the diverse obstacles students encounter throughout their learning journey. Solving the problems encountered by students poses a significant challenge for traditional deep learning models, as it requires not only a broad spectrum of subject knowledge but also the ability to understand what constitutes a student’s individual difficulties. It’s challenging for traditional machine learning models, as they lack the capacity to comprehend students’ personalized needs. Recently, the emergence of large language models (LLMs) offers the possibility for resolving this issue by comprehending individual requests. Although LLMs have been successful in various fields, creating an LLM-based education system is still challenging for the wide range of educational skills required. This paper reviews the recently emerged LLM researches related to educational capabilities, including mathematics, writing, programming, reasoning, and knowledge-based question answering, with the aim to explore their potential in constructing the next-generation intelligent education system. Specifically, for each capability, we focus on investigating two aspects. Firstly, we examine the current state of LLMs regarding this capability: how advanced they have become, whether they surpass human abilities, and what deficiencies might exist. Secondly, we evaluate whether the development methods for LLMs in this area are generalizable—that is, whether these methods can be applied to construct a comprehensive educational supermodel with strengths across various capabilities, rather than being effective in only a singular aspect. Based on the current development status, we further outline two approaches for an LLM-based education system: a unified approach and a mixture-of-expert (MoE) approach. Finally, we explore the challenges and future directions, providing new research opportunities and perspectives on adapting LLMs for education. footnotetext: $\dagger$ Weinan Zhang, Yong Yu and Wei Xia are the co-corresponding authors.
Large Language Models, Educational Data Mining
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systemsccs: Applied computing Education
1. Introduction
@@
Education plays a vital role in shaping individuals’ futures as it forms the foundation for providing people with knowledge, skills, and critical thinking abilities. Conventional education systems heavily rely on teachers for imparting knowledge to students, which place a significant demand on educational resources. However, the advent of online education has substantially lowered the cost of accessing these educational materials. Many people are conveniently acquiring knowledge through online courses and exercises.
Much effort has been made to achieve personalized learning in online-education (Abdelrahman et al., 2023; Li et al., 2023b; Gong et al., 2020). Most of these methods are based on predicting students’ knowledge states or recommending personalized learning resources using neural networks, based on sequences of student behaviors. However, this approach achieves only a coarse level of personalization. Even when students receive recommended resources, the specific difficulties they encounter in their learning process can remain unresolved. These difficulties may vary from student to student; for example, different students might struggle with different aspects of the same problem, such as misconceptions in understanding key concepts or difficulties in the reasoning process. These issues require detailed descriptions from each student for educators to understand. However, current online education systems face the issue of not being able to interact with students in real-time like a teacher. Presently, online education platforms typically offer static course videos and exercises, leaving students unable to ask questions or seek dynamic solutions to their specific issues. Therefore, developing a teaching assistant model capable of addressing students’ individual concerns represents a crucial step forward in advancing online education.
@@
The emergence of large language models (LLMs) instills optimism in creating an intelligent education system. Since the launch of ChatGPT, LLMs have shown exceptional ability in understanding human knowledge and have been widely applied in various professional fields, including recommendation systems (Lin et al., 2023), healthcare (Liu et al., 2023f), economics (Li et al., 2023a), and others. By introducing expansive world knowledge, considerable reasoning capabilities, and the ability to understand human language, LLMs have the potential to introduce new forms of interactions and methodologies across these fields. However, the challenge intensifies when constructing an LLM-based education system. To solve the specific problems encountered by students, it’s not only necessary for LLMs to understand the precise issues faced by the students but also essential for them to possess specialized educational-related knowledge and skills. Only then they would be able to address problems that the students themselves might not be able to resolve.
@@
The act of LLMs answering students’ questions can be considered as a process that involves using multiple educational capabilities. For instance, as depicted in Figure 1, the question posed by the student requires the education assistant to possess both mathematics and programming skills simultaneously. Therefore, summarizing the methods of developing LLMs’ education-related capabilities is meaningful for building the next-generation intelligent education system. In this paper, we study LLMs from the perspective of these education-related capabilities and explore the potential of an LLM-based education system, aiming to provide valuable insights into the areas that need further improvement. We present our primary investigation in Figure 2. Our specific focus centers on five distinct capabilities of LLMs:
Figure 1. An example of LLM-based education systems integrating multiple abilities to solve student problems.
@@
  • Mathematics: We primarily summarize the current development of LLMs in solving mathematical problems of varying difficulties and types, encompassing basic arithmetic operations, challenging problems, multi-modal problems, and math proof.
@@
  • Programming: In accordance with human programming conventions, we divide LLMs’ programming process into two stages: code writing and code refinement. We review the researches in this area and summarize the remaining problems for LLMs in coding.
@@
These five capabilities form the foundation of an LLM-based education system. Building upon this foundation, we propose two potential approaches for forming an LLM-based education system. One approach involves training a comprehensive language model with multiple capabilities, while the other is based on an LLM controller in a mixture-of-experts framework.
The rest of this paper is organized as follows. In section 2, we briefly introduce the education tasks and educational LLMs. In section 3, we summarize the current development status of five foundational abilities related to education, followed by discussion of the development trend of education-related LLM capabilities. In section 4, we investigate the performance of well-known LLMs in education-related capabilities. In section 5, we introduce the potential approaches for organizing an LLM-based education system. Finally, in section 6 and section 7, we highlight the challenges and future directions for designing an LLM-based education system and concludes this survey.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, align=left, minimum width=2.5em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.6em,font=, where level=2text width=5.3em,font=, where level=3text width=9.7em,font=, [Foundational
@@
Capabilities,text width=6.3em [Mathematics, text width=5.5em [Basic Arithmetic Problems, text width=10em [Goat (Liu and Low, 2023), GENBERT (Geva et al., 2020),leaf, text width=11.2em] ] [Challenging Mathematics Problems,text width=13.3em [WizardMath (Luo et al., 2023a), MeatMath (Yu et al., 2023),leaf, text width=13.5em] ] [Multimodal Problems,text width=8.3em [Mathvista (Lu et al., 2023), Unigeo (Chen et al., 2022a),leaf, text width=11.5em] ] [Mathematical Proof,text width=8em [LeanDojo (Yang et al., 2023b), LEGO-Prover (Xin et al., 2023),leaf, text width=13.5em] ] ] [Writing,text width=4em [Text Summarization,text width=8.0em [BRIO (Liu et al., 2022b), PROM (Ma et al., 2023a),leaf, text width=10em] ] [Grammatical Error Correction,text width=11.5em [CoEdit (Liu et al., 2022b), GrammarGPT (Fan et al., 2023),leaf, text width=12em] ] ] [Programming,text width=6em [Code Writing, text width=5.2em [CodeLlama (Rozière et al., 2023), WizardCoder (Luo et al., 2023b),leaf, text width=13.5em] ] [Code Refinement, text width=7.2em [Codet (Chen et al., 2022b), SEIDR (Liventsev et al., 2023),leaf, text width=10em] ] ] [Reasoning, text width= 5em [Supervised Fine-tuning,text width=8.6em [CAGE (Rajani et al., 2019), Scratchpad (Nye et al., 2021),leaf, text width=11.3em] ] [Prompt Engineering,text width=7.8em [CoT (Wei et al., 2022), SelfConsistency (Wang et al., 2022),leaf, text width=12em] ] [Hybrid Methodologies,text width=8.3em [FineTuneCoT (Ho et al., 2022), STaR (Zelikman et al., 2022),leaf] ] ] [Knowledge-based QA , text width= 8.3em [Open-domain QA , text width= 7.5em [Retro (Borgeaud et al., 2022), FRESHPROMPT (Vu et al., 2023),leaf, text width=12em] ] [Domain-specific QA,text width=7.5em [Reta-llm (Liu et al., 2023d), LeanContext (Arefeen et al., 2023),leaf]] ] ] ] ]
Figure 2. Summary of LLM’s education-related foundational capabilities.
2. Background
In this section, we first discuss the educational tasks and introduce of LLMs’ roles in smart education. Then, we introduce the current development of educational large language models and compare our survey with previous works.
2.1. Educational Tasks
Artificial Intelligence can significantly boost the development of online education. Developing an intelligent education system involves tackling a wide array of tasks, which can be broadly categorized into two types. The first type centers around solving personalized, knowledge-based questions from students(Zhou et al., 2023a; Yu et al., 2023; Yasunaga et al., 2021). This category aims to assist students in resolving their queries during the learning process, such as clarifying misunderstandings about a particular concept(Lazaridou et al., 2022), solving a math problem(Wang et al., 2017; Gou et al., 2023; Xiong et al., 2023), or writing a piece of code to address a specific issue(Chen et al., 2022b; Roziere et al., 2023; Zhang et al., 2023a). The second type focuses on aiding students with their learning planning, such as mapping out learning paths(Li et al., 2023b; Chen et al., 2023), knowledge tracing(Abdelrahman et al., 2023; Piech et al., 2015; Corbett and Anderson, 1994), and computerized adaptive testing(Ghosh and Lan, 2021; Meijer and Nering, 1999; Thompson and Weiss, 2019). These tasks are designed to support the learning process from a broader perspective, instead of addressing specific learning challenges students face.
In this paper, we mainly survey the former category which mainly develops the abilities of LLMs answering students’ specific questions requiring specific skills, due to that it’s the main scenario that LLMs could contribute more. The former scenario is where LLMs currently play a significant role, for two main reasons: 1) From the perspective of task characteristics, tasks like designing learning paths and tracking knowledge, though also guiding student learning, are mainly based on students’ learning sequences. The reasoning process mostly happens in the background, with relatively less need for dialogue. 2) From the viewpoint of the characteristics of LLMs, the advantage of language models over traditional recommendation models lies in their extensive world knowledge, conversational abilities, and logical reasoning capacities. These capabilities are crucial for addressing the personalized and subject-specific questions that students encounter. These questions are often complex and personalized, requiring dialogue for effective understanding—a feature not possessed by previous deep learning models. However, for issues like learning path planning and knowledge tracking, deep learning models, through training, can handle them well (Piech et al., 2015; Li et al., 2023b; Ghosh and Lan, 2021).
2.2. Educational Large Language Models
Currently, many online education companies have launched their own large educational models(iFLYTEK, 2024; Group, 2024). The iFLYTEK Spark(iFLYTEK, 2024), introduced by iFLYTEK Company, boasts capabilities in multimodal interaction, coding, text generation, solving math problems, and knowledge Q&\&&A. The introduction of LLMs could evidently enhance the efficiency of learners (Kazemitabaar et al., 2024). MathGPT(Group, 2024), developed by TAL Education Group, is an LLM specializing in math problem-solving and lecturing. From the development patterns of these companies, it is evident that the primary application scenario for these large models is to address specific knowledge-based questions from students. This approach benefits the industry in two main ways: 1) It allows for more interaction with students. Answering specific questions about subjects or knowledge points involves more detailed participation in the students’ learning process compared to learning path planning and knowledge tracking, thereby increasing the time students spend on the platform. 2) Solving student questions makes the model appear more intelligent. Planning learning paths and assessing students can be achieved with traditional models, but understanding and solving students’ problems through dialogue is only possible with large models, making the platform’s products more intelligent. In summary, because they better retain student users and make their products more intelligent, the main trend for platforms developing educational large models is to enable them to solve specific student questions. Of course, issues like learning path planning and knowledge tracking are also very important, but the transformation in the era of large models requires further exploration.
2.3. Related Surveys
LLMs hold vast potential in the field of education. There are already several surveys about LLMs in education, while our work is different from them. Gan et al. (2023) explored various roles LLMs can play in the education process. It focuses on analyzing the roles undertaken by LLMs from the perspective of different application scenarios, such as learning support tools, personalized learning experiences, content creation and generation, language learning and teaching, cross-lingual communication, and translation. The basic capabilities of LLMs are not discussed in it.  Kasneci et al. (2023) explores the benefits and challenges of LLMs in education from both students’ and teachers’ perspectives, highlighting the potential of LLMs in research, writing, and problem-solving tasks, as well as their ability to provide domain-specific language skills for specialized learning.
Other than that, AL-Smadi (2023) primarily explored the use of generative AI models in education, focusing on their application as teaching aids, the generation of personalized learning materials, and the assessment of student learning outcomes. It primarily assesses ChatGPT’s performance on tasks such as instructional design from an educational perspective, in a qualitative rather than quantitative manner.  Meyer et al. (2023) is an editorial on the opportunities and challenges of LLMs in academia, analyzing the potential impacts and risks of LLMs from the perspectives of academic writing, education, and programming education. They primarily discussed the role of LLMs in education from a pedagogical perspective, whereas our work leans towards analyzing the current capabilities of LLMs in solving subject-specific questions in the educational process from a technological standpoint, offering ideas and insights for creating an LLM capable of addressing problems across various subjects. Wang et al. (2024) summarized the development of LLMs in education from the perspective of data and technology, discussing the current applications of LLMs in tasks such as study assisting, teach assisting, and adaptive learning. It is important to note that while the article mentions some technical methods for developing LLMs’ problem-solving abilities in study assisting, the content covered is not systematic. The technical development for solving problems for students didn’t take up much space of this article. Moreover, the disciplines explored in the article are not foundational but include advanced subjects such as Medicine and Finance. In contrast, our work primarily focuses on discussing the development of LLMs’ fundamental capabilities in assisting students with solving various problems.
While previous surveys have provided ample discussion on the potential applications of LLMs in education, they exhibit two main shortcomings: 1) Their exploration of LLMs’ applications in education often spans a wide range of topics, including designing learning paths, assisting teachers, and planning curricula. As discussed in Section  2.1, although these capabilities are important, they do not represent the primary direction of current practical applications of educational LLMs. Our work, in contrast, primarily focuses on the ability of LLMs to answer subject-specific questions. 2) They have not analyzed the development of LLMs’ educational capabilities from a technological perspective. Discussing the evolution of LLMs’ educational abilities from a technical standpoint is crucial for building a general LLM-based intelligent education system. Different from them, we review the evolution of LLMs from the perspective of education-related capabilities. We summarize the techniques to promote LLMs in these capabilities. Additionally, we provide foresight into constructing feasible frameworks for LLM-based education systems. Our work emphasizes a comprehensive understanding of LLMs’ educational competencies and explores frameworks that can effectively integrate these abilities into the educational landscape.
3. Foundational Capabilities
3.1. Mathematics
@@
Mathematics demands reasoning of complex information, making it one of the disciplines that place the highest premium on the cognitive abilities. There is significant academic interest in enhancing the mathematical capabilities of LLMs (Lu et al., 2022b). In the pursuit of creating an education system based on LLMs, the goal is to equip it with the capability to tackle a wide range of mathematical problems. These problems may encompass basic numerical calculations, complex logical reasoning, or challenges that require the integration of information from multiple modalities. In this section, we summarize the developments of the mathematical capabilities of LLMs, focusing on four aspects: fundamental numerical computations, complex reasoning, the handling of multi-modal problem-solving, and mathematical proof.
3.1.1. Basic Arithmetic Problems
@@
Recently, significant scholarly attention has focused on augmenting LLMs’ proficiency in this domain (Patel et al., 2021; Wang et al., 2017; Zhang et al., 2023b; Geva et al., 2020). In human learning mathematics, foundational mathematical operations serve as the basis for addressing more advanced problems. Given the robust comprehension of human language and notable textual reasoning abilities exhibited by LLMs, it is natural for many to assume that LLMs should effortlessly handle fundamental mathematical problems. However, the reality proves otherwise. Yuan et al. (2023) pointed out that ChatGPT and GPT-4 (Achiam et al., 2023) perform well in addition and subtraction operations, but their accuracy decreases when dealing with multiplication involving larger numerical values. This limitation arises because the LLMs do not access a calculator during the computations. More importantly, when LLMs solve computational problems, their internal logic does not perform actual calculations. Instead, they generate text to predict each digit of the answer step by step. This approach means that as the number of digits in the answer increases, the probability of making an error grows cumulatively.
This problem is not unsolvable. Yang et al. (2023a) and Liu and Low (2023) proposed to fine-tune LLMs on high-quality datasets and found that even small language models could avoid making mistakes on multi-digit problems. Lee et al. (2023) found that even small transformers could solve arithmetic problems with high accuracy as long as trained on data with detailed calculation process. All in all, fine-tuning on high quality data is one workable solution. However, this approach may only work when a specialized arithmetic model need to be build while not applicable for building general LLMs since it is not feasible to fine-tune each LLM individually. How to prevent such issues during the pre-training phase of LLMs remains an open question. At the current stage, a simple and feasible solution is to have the LLMs outsource arithmetic problems to a calculator, which can ensure the accuracy of the calculations (Schick et al., 2024).
3.1.2. Challenging Mathematics Problems